# Enhancing selectivity using Wasserstein distance based reweighing

### v1

@@ Abstract
On the theoretical side, we prove that when the metric entropy of the input data-sets is bounded, our greedy algorithm outputs a close to optimal reweighing, i.e., the two invariant distributions of network weights will be provably close in total variation distance. Moreover, the algorithm is simple and scalable, and we prove bounds on the efficiency of the algorithm as well.
@@ 1 Introduction
Deep learning has found applications in diverse areas ranging from organic chemistry to computer generated art. Its applicability is limited by the availability of large amounts of labeled training data. A typical all-to-all neural net of width $n$ (say $n=10000$) and depth $d$ (say $d=10$) has $\Theta(n^{2}d)$ ($\simeq 10^{8}$) weight parameters, and although many neural nets work well despite being somewhat over-parameterized, they still need the number of training examples to be of a similar or only slightly smaller order of magnitude.
This reliance on large amounts of training data leads to difficulties in training when we have multiple objectives. For example, suppose we separately gather training data for a neural net to perform classification tasks into a set of classes A using labeled data-set $\mathcal{S}$, and into a set of classes B using labeled data-set $\mathcal{T}$. Then, unless $\mathcal{T}$ and $\mathcal{S}$ have many data-points in common, we will not have enough examples to train a neural net for the classification into classes of $A\times B$. In this paper, we address this problem by designing a scalable algorithm that reweighs dataset $\mathcal{S}$ (using $\mathcal{T}$), so that training on the reweighted $\mathcal{S}$ leads to network that is close to one obtained by training on $\mathcal{T}$. Moreover, we apply it to a drug discovery application and obtain wet-lab verified results.
@@
Our main contribution is theoretical. Algorithm 1 deliberately introduces distribution skew and reweighs the labeled training data-set $\mathcal{S}$ using the data-set $\mathcal{T}$ so that if we train a neural network on the reweighed $\mathcal{S}$ for a long enough period of time then its weights will be "tilted" so that the classification error on classes $B$ in $\mathcal{T}$ will also be reduced. The amount of reduction is determined by the choice of tilt parameter $\alpha$ in Algorithm 1. Theorems 5.45.135.15, and 5.16 formally show correctness and efficiency of Algorithm 1. In particular, Theorem 5.4 proves correctness and also justifies the choice of Wasserstein metric in Algorithm 1. Theorems 5.135.15 and 5.16 provide efficiency guarantees for Algorithm 1.
@@
It has been known since the 1980s that greedy algorithms, like Algorithm 1, have poor approximation guarantees for computing Wasserstein distance [20]. Therefore, we need to assume and exploit some property of about our input instances to get around the lower bound in [20]. This critical property turn out to be: input instances must admit a small size covering. This small covering assumption is used in both the random sampling and the greedy reweighing steps of our algorithm. The connection to random sampling is not surprising and it comes from known techniques – the union bound in the large deviations proof. However, the connection between small coverings and greedy matching algorithms (at its core computing Wasserstein distance is equivalent to computing minimum weight matchings) is somewhat surprising, since we are not aware of results obtaining sharper guarantees on approximate minimum weight matching, based on the covering properties of the input data-set.
@@
Organization: In Section 2, we describe an example application to drug discovery. More theoretically inclined readers may skip this section and any references to it. In Section 3, we discuss prior work from the areas of machine learning theory, algorithms and computational drug discovery, relevant to our paper. In Section 4, we present Algorithm 1 and provide a technical overview of our paper – how our various theoretical results fit together. Section 5 presents the theorem statements. In particular, Theorem 5.4 explains why the Wasserstein metric is an intuitive and appropriate choice of metric for Algorithm 1; Theorems 5.135.15 and 5.16 show that the greedy random sampling based algorithm can compute the minimum weight bipartite matching, and hence Wasserstein distance, in near linear time; and they provide an upper-bound on the approximation error under our low metric entropy assumption. Thus showing that Algorithm 1 is scalable. Finally, the supplement contains missing proofs.
2 Example application: drug discovery
Multicriteria optimization problems involving data-sets that admit small coverings arise naturally in drug discovery (see Section 4 for details). Therefore, as a concrete motivating example, we illustrate an application of Algorithm 1 to a toy problem in this area.
In drug discovery, typically, one wants to isolate small molecules (inhibitors) that bind strongly to a given enzyme, but often we want to exclude small molecules that bind to another similar enzyme. For example, MNK1 and MNK2 are two structurally similar kinases (a kinase is an enzyme for phosphorylation or de-phosphorylation of proteins). We want to identify small molecules that bind strongly to MNK2 (MNK2 hits), but we also prefer that the identified small molecules not bind to MNK1 (MNK1 non-hits). In other words, we want to isolate molecules that are selective for MNK2 over MNK1.
In the in-silico experiments, we were able to increase the percentage of MNK1 non-hits in our set of top predicted MNK2 hits – the selectivity – from 54% to 95% on holdout data, using the reweighing procedure in Algorithm 1. We used a relatively small training set of about $250$K small molecules in total; labeled as MNK1 non-binders, and MNK2 binders and non-binders; and a small holdout set of $7$K small molecules that consists of molecules which are labeled as: MNK2 hits (binders), and MNK1 hits (binders) or MNK1 non-hits (non-binders). In Figure 1, for the neural network models with and without reweighing, we plot the cumulative number of MNK1 non-hits on the $y$-axis; and on the $x$-axis any given point, say $k$, represents the top $k$ predicted MNK2 hits from the examples in the holdout set.
While we can not make our training data-sets and code public for proprietary reasons, we were able to experimentally (in wet-lab) verify that two out of the top fifty (actually 43, since 7 out of 50 molecules could not be synthesized and tested) predicted selective small molecules, obtained by running our neural network model on the Enamine 1.9B molecules catalog (https://enamine.net), were indeed selective for MNK2 over MNK1.111The compounds are Z1918489591 and Z5890616727 in the enamine catalog. That is a success rate of roughly 5% on this admittedly small sample set. We do note that the results are from a single point concentration assay and can be noisy.
We are not aware of other such multi-target prediction results in DNA encoded library (DEL) space (see [22] for background), where one simultaneously predicts hits/non-hits against two or more proteins. However, the success rates for single target experiments with traditional high-throughput screening is $\sim 1\%$ (see for example the discussion in [17]) and it is generally accepted that multi-target prediction is a harder problem.
More importantly, the two predicted and assay tested molecules provide a degree of verification for our experimental application (which has been the motivation, but is not the focus of this paper).
@@ 3 Related work
The question of learning with differing test and train distributions has been well investigated in the machine learning community under different names: ditribution shift and covariate shift, see for example the book [19], the papers [24, 21, 10, 3, 13] and [7], to name just a few. The question is also relevant to our paper since our algorithm can be used to reweigh the train data-set to bring the post training neural network weights closer to what they would have been, had we trained based on the test distribution. However, prior results rely on estimating the train and test distributions. For example, parameter estimation of the densities followed by change of variable using the Jacobian. Assuming a logarithmic number of features, that leads to a $\tilde{O}(n^{2})$ algorithm for distribution skew correction ($n$ being the training and test data-set size) – much more efficient than Wasserstein distance computation that requires solving a $\Theta(n^{2})$ sized linear program. That explains why the Wasserstein distance based ideas are less explored in this context, so far. Distribution shift correction has also found new applications in domain adaptation literature. However, the only prior theoretical works involving Wasserstein distance computation that we found in this area were [8] and [16], which focus on exact solution of the Wasserstein distance problem.
In a high dimensional feature space, the density estimation can be inaccurate since the number of samples required increases exponentially in number of features for any formal guarantee (as can be seen from large deviation bounds [9]). Moreover, if we are only interested in partial tilting of one distribution towards another, as in Algorithm 1 (where $\alpha$ controls the amount of tilt), then it is reasonable to look for approximate but efficient computation of Wasserstein distance. That is what we do in this paper using Algorithm 1, which is a $\tilde{O}(n^{2})$ time algorithm as well. We now have the added advantage of a provable upper bound on the test-train prediction error, before and after tilting using Algorithm 1, from Theorem 5.4.
@@
The problem of efficient Wasserstein distance computation has also received much attention in the algorithms community. The paper [23] studies the equivalence between Wasserstein distance computation and matching algorithms in the metric space setting. Efficient matching algorithms have been well studied in literature for five decades. The optimal algorithm for computing weighted matchings is due to Gabow and Tarjan [11] and runs in time $O(m\sqrt{n})$, where $m$ is the number of edges and $n$ the number of vertices in the graph. Since then more sophisticated algorithms have been designed, see for example  [25, 14, 23] and [1] to name a few. However, under our assumptions even the simple greedy algorithm performs remarkably well, and it scales efficiently for large training data-sets.
@@
Reingold and Tarjan [20] showed that the greedy algorithm has an abysmal approximation ratio of $n^{\log_{2}3/2}$ for bipartite graphs. In this paper, we show in Theorem 5.13 that the approximation ratio of the greedy algorithm is much better under our bounded metric entropy assumption than the lower bound in [20]. Hence, an assumption about a covering property of the input leads to more optimal matchings – a somewhat surprising algorithmic result that may be of independent interest.
@@
Finally, the idea of using deep learning for drug discovery has gained popularity in pharmaceutical research over the last few years, especially given the amount of data now available [18]. The paper [17] shows that neural nets can be trained on DNA encoded chemical libraries to identify new small molecules that bind to a given protein target. It is particularly relevant to this work, as we build upon that. Our work extends their work by allowing us to select molecules that bind to one protein target and not to another. Other papers in this rapidly growing area include [26, 15] and [12].
@@ 4 Problem statement and overview of results
Suppose we are given two labeled training data sets, say $\mathcal{S}$ and $\mathcal{T}$ for two different classification tasks. Moreover, let’s assume that the points in the data-sets are weighted according to two different probability distributions, say $\mathbb{P}_{\mathcal{S}}$ and $\mathbb{P}_{\mathcal{T}}$ respectively. Our goal is to reweigh $\mathcal{S}$, i.e., "tilt" $\mathbb{P}_{\mathcal{S}}$ towards $\mathbb{P}_{\mathcal{T}}$, and train a neural net classifier so that the limiting distribution of network weights is closer to the one that would be obtained from training using $\mathbb{P}_{\mathcal{T}}$. We assume labels of $\mathcal{S}$ are known, and the labels of $\mathcal{T}$ may be unknown or they may be known but $|\mathcal{T}\cap\mathcal{S}|$ may be small. Our reweighing algorithm handles both cases. Furthermore, our reweighing procedure (Algorithm 1) works efficiently on very large data-sets, with provable guarantees.
@@
For our drug discovery example, the set $\mathcal{T}$ consists of a subset of small molecules [^2] As an aside, each small molecule is usually mapped to a 2K character long binary string (fingerprint) of features. Thus, in this context, one may think of the underlying space of small molecules as a subset of the boolean hypercube in dimension 2K. labeled non-binders (non-hits) for the protein MNK1, and the full labeled training set consists of molecules that are binders and non-binders for the protein MNK2, while the set $\mathcal{S}$ is not the full training set but just the set of binders to MNK2333Note that we have the labeled set of non-binders to MNK2 in our training examples as well, but they remain unaffected by the reweighing, though they are used in training as well. . Here the labels of the molecules in $\mathcal{T}$ are known but not necessarily on the same molecules as $\mathcal{S}$. $\mathbb{P}_{\mathcal{S}}$ and $\mathbb{P}_{\mathcal{T}}$ may be assumed to be uniform distributions supported on $\mathcal{S}$ and $\mathcal{T}$ respectively. Given a new small molecule, one now wants to compute the likelihood that it is a binder for MNK2 and a non-binder for MNK1. Such models can allow us to make predictions on large commercially available catalogs and enrich compounds that have high likelihood to bind to MNK2 but not MNK1. This formulates the drug discovery application in our formal notation.
The rest of the paper concentrates on providing a theoretical explanation for why Algorithm 1 should work as intended and scale well in general; beyond the specific experiments with MNK1-MNK2.

### v2

@@ Abstract
On the theoretical side, we prove that when the metric entropy of the input datasets is bounded, our greedy algorithm outputs a close to optimal reweighing, i.e., the two invariant distributions of network weights will be provably close in total variation distance. Moreover, the algorithm is simple and scalable, and we prove bounds on the efficiency of the algorithm as well.
@@ 1 Introduction
Deep learning has found applications in diverse areas ranging from organic chemistry to computer generated art. Its applicability is limited by the availability of large amounts of labeled training data. This reliance on large amounts of training data can lead to difficulties in training. For example, suppose we separately gather training datasets $\mathcal{S}$ and $\mathcal{T}$, that consist of color images of cats and dogs, for a neural net to perform two different classification tasks: (1) classify cats vs dogs and (2) classify red vs blue. Now, if we want a new classifier that uses $\mathcal{S}$ to more accurately distinguish red cats vs other colored cats, then the usual approach would be to train another neural net on labeled samples of red cats vs others. If the category red cats represents disproportionately few labeled samples in $\mathcal{S}$ compared to $\mathcal{T}$, then we are potentially out of luck, unless we gather more labeled data. In general, this may be unavoidable, but when the underlying ground space for the two sampled datasets is the same, then we may be able to use information from $\mathcal{T}$ in $\mathcal{S}$, especially if nearby objects are likely to be mapped to the same class. In this paper, we design a scalable algorithm (see Algorithm 1) that reweighs points of $\mathcal{S}$ (using $\mathcal{T}$), so that the limiting distribution of post-training network weights using reweighted $\mathcal{S}$ is provably close to what would be obtained by training on $\mathcal{T}$; or more generally, some user specified weighted mixture of $\mathcal{S}$ and $\mathcal{T}$ (see theorems in Section 3.3). We also demonstrate the effectiveness of our reweighting algorithm in a practical application: to discover selective small molecule kinase inhibitors, and we obtain in-silico (see Figure 1), as well as assay verified results (see Figure 2).
@@
Our main contributions are theoretical. Algorithm 1 reweighs the labeled training data-set $\mathcal{S}$ using the data-set $\mathcal{T}$ so that if we train a neural network on the reweighed $\mathcal{S}$ for a long enough period of time then the network weights will be ”tilted” so that the classification error with respect to $\mathbb{P}_{\mathcal{T}}$ will be reduced. The amount of reduction is determined by the choice of tilt parameter $\alpha$ in Algorithm 1. Theorems 3.23.33.4, and 3.5 formally show correctness and efficiency of Algorithm 1. In particular,
  • Theorem 3.2 justifies the choice of Wasserstein metric in Algorithm 1,
@@
  • Theorem 3.4 shows that greedy algorithm has a sub-logarithmic approximation guarantee for minimum weight metric bipartite matching on the boolean hypercube, when the underlying dataset has small metric entropy (roughly equivalent to existence of a small covering). In Reingold and Tarjan (1981), the authors showed that the greedy algorithm has poor approximation guarantees for computing minimum weight metric bipartite matchings. [^1] At its core, computing Wasserstein distance is equivalent to computing minimum weight metric bipartite matchings, see for example Sharathkumar and Agarwal (2012). Therefore, we need to assume and exploit some property of our input instances to get around the lower bound in Reingold and Tarjan (1981). Assuming that input instances have small metric entropy turns out to be sufficient. This connection between small coverings and greedy matching algorithms is somewhat surprising, since we are not aware of results obtaining sharper guarantees on approximate minimum weight matching, based on the covering properties of the input data-set.
  • Theorem 3.4 and Theorem 3.3 provide (multiplicative) approximation guarantees for the greedy and randomized greedy algorithm for approximate $1$-Wasserstein distance computation.
@@
Organization: In Section 2, we discuss prior work from the areas of machine learning theory, algorithms and computational drug discovery, relevant to our paper. In Section 3, we present Algorithm 1 and provide a technical overview of our paper – how our various theoretical results fit together. In particular, Theorem 3.2 explains why the Wasserstein metric is an intuitive and appropriate choice of metric for Algorithm 1; Theorems 3.43.5 and 3.3 show that the greedy random sampling based algorithm can compute the minimum weight bipartite matching, and hence Wasserstein distance, in near linear time; and they provide an upper-bound on the approximation error under our low metric entropy assumption. Thus showing that Algorithm 1 is scalable. In Section 4, we describe an example application to drug discovery. Finally, the supplement describes the formal setup, theorem statements, proofs; and provides further figures and details about our drug discovery application.
@@ 2 Related work
The related question of learning with differing test and train distributions has been well investigated in the machine learning community under various names, including domain shift and distribution shift, see for example the books Quionero-Candela et al. (2009) and Redko et al. (2019), the papers Shimodaira (2000); Rosenbaum and Rubin (1983); Dudik et al. (2005); Bickel and Scheffer (2007); Huang et al. (2007); Ben-David et al. (2010) and Cortes et al. (2008), to name just a few. The question is also relevant to our paper since our algorithm can be used to reweigh the train data-set to bring the post training neural network weights closer to what they would have been, had we trained based on the test distribution. Some prior results rely on estimating the train and test distributions. For example, parameter estimation of the densities followed by change of variable using the Jacobian. Assuming a logarithmic number of features, that leads to a $\tilde{O}(n^{2})$ algorithm for distribution skew correction ($n$ being the training and test data-set size) – much more efficient than Wasserstein distance computation that requires solving a $\Theta(n^{2})$ sized linear program. However, in high dimensional feature spaces, the number of samples required increases exponentially in number of features, for any formal guarantee for density estimation (as can be seen from large deviation bounds Dembo and Zeitouni (2010)). Moreover, if we are only interested in partial tilting of one distribution towards another, then it is reasonable to look for approximate but efficient computation of Wasserstein distance. That is what we do in this paper using Algorithm 1. Algorithms for distribution shift correction are applied to domain adaptation as well, as the two problems are very similar. The amount of literature in domain shift and distribution shift is vast. However, the only prior theoretical works involving Wasserstein distance computation that we found in this area were Courty et al. (2017) and Le et al. (2021), which focus on exact solution of the Wasserstein distance problem.
@@
The problem of efficient Wasserstein distance computation has also received much attention in the algorithms community. The paper Sharathkumar and Agarwal (2012) studies the equivalence between Wasserstein distance computation and matching algorithms in the metric space setting. Efficient matching algorithms have been well studied in literature for five decades. The optimal algorithm for computing weighted matchings is due to Gabow and Tarjan Gabow and Tarjan (1991) and runs in time $O(m\sqrt{n})$, where $m$ is the number of edges and $n$ the number of vertices in the graph. Since then more sophisticated algorithms have been designed, see for example  Vaidya (1989); Imielinska and Kalantari (1993); Sharathkumar and Agarwal (2012) and Andoni et al. (2009) to name a few. However, under our assumptions even the simple greedy algorithm performs remarkably well, and it scales efficiently for large training data-sets.
@@
Note that Reingold and Tarjan (1981) showed that the greedy algorithm has an abysmal approximation ratio of $n^{\log_{2}3/2}$ for bipartite graphs. In this paper, we show in Theorem 3.4 that the approximation ratio of the greedy algorithm is much better under our bounded metric entropy assumption than the lower bound in Reingold and Tarjan (1981). Hence, an assumption about a covering property of the input leads to more optimal matchings – a somewhat surprising algorithmic result that may be of independent interest.
In the context of approximate Wasserstein distance computation, we are also aware of the Sinkhorn algorithm. Sinkhorn distance computation (gradient ascent) can take about a second for 2K points, [^2] For a comparison of various Sinkhorn distance algorithms, see Figure 4 in Cuturi (2013). which is likely going to be slower than most implementations of the greedy matching algorithm (sorting). Sinkhorn accuracy can be traded-off with computation time. However, as accuracy is decreased, the issue of worst case approximation guarantee becomes relevant. We are aware of additive approximation guarantees for Sinkhorn (see Theorem 1 in Genevay et al. (2019)) but not multiplicative ones. Directly interpreting their result, in our setting (the $d$ dimensional hypercube), their additive guarantee is: $\Theta(d\log d)$, while the diameter of the space is $d$. Hence the trivial upper-bound on the Wasserstein distance is $O(d)$, which makes their guarantee not useful for us. We are not aware of worst case multiplicative approximation guarantees for the Sinkhorn algorithm. We show worst case multiplicative approximation guarantees for the greedy algorithm in Theorem 3.3. The approximation factor can be informally summarized as: $O(d^{c})$ for some constant $c<1$. The exact $c$ achieved depends upon the size and radii of the balls in the optimal covering, which may be unknown, and an upper bound is used in the statement. Obtaining guarantees for the Sinkhorn algorithm, under the small covering assumption here, is an interesting open problem.
One can use different objectives in the problem setting, including reweighing to minimize the generalization error. Our objective is to bring the limiting distribution of neural net weight parameters closer. [^3] It is worth noting here that while it is possible to construct examples (like the exponential function) where small weight perturbations lead to large perturbations in output; such (non-robust) neural nets are less likely to be useful. For example, robustness (to weight perturbations) is desirable for model compressibility (see the discussion in section ”Related works” of Tsai et al. (2021)). Moreover, generalization error of SGD is upper-bounded by the sensitivity (opposite of robustness) of the square of the gradient under weight perturbations (see Theorem 1 in Neu et al. (2021)). Thus the distribution of the invariant measures of weights being close is a reasonable metric for neural net weight parameters; as robustness to weight perturbations is desirable for neural nets, for reasons above. In Neu et al. (2021), they explore upper-bounds on the generalization error in terms of the statistical properties of stochastic gradient descent (SGD). [^4] Their bound is in terms of the variance of the gradients around the local minima. We suspect that using it together with the bound in Theorem 3.2 in this paper, would lead to a similar bound on the generalization error (in the limit as $t\to\infty$ and step-size is small.) but now as an expectation over network weights, where the expectation is with respect to the invariant measure of the limiting SDE corresponding to the SGD. We do not know of any bounds in the reverse direction, but it seems to be a harder problem to judge how far apart are the network weights given the generalization error is small or large. Discrepancy minimization for reducing generalization error has been explored recently in the context of a very similar problem to ours (see Awasthi et al. (2024), and the references therein). Such approaches usually solve a convex program, while we solve a nearest neighbor problem. The latter is more tractable for large datasets. Moreover, such approaches often do not give an approximation guarantee, but we do. Furthermore, Awasthi et al. (2024) does not take into account the training algorithm, i.e., that we used stochastic gradient descent (SGD) to train our model. The choice of SGD for training motivated our choice of Wasserstein distance in Algorithm 1 (see Theorem 3.2).
@@
Finally, the idea of using deep learning for drug discovery has gained popularity in pharmaceutical research over the last few years, especially given the amount of data now available Mullard (2016). The paper McCloskey et al. (2020) shows that neural nets can be trained on DNA encoded chemical libraries to identify new small molecules that bind to a given protein target. It is particularly relevant to this work, as we build upon that. Our work extends their work by allowing us to select molecules that bind to one protein target and not to another. Other papers in this rapidly growing area include Zhou et al. (2019); Kearnes et al. (2016) and Gilmer et al. (2017).
@@ 3 Problem statement and overview of results
Suppose we are given two training data sets, say $\mathcal{S}$ and $\mathcal{T}$, consisting of $\Theta(n)$ points each, for two different classification tasks. Assume that the labels of $\mathcal{S}$ are known, and the labels of $\mathcal{T}$ are unknown. [^5] Or the labels of $\mathcal{T}$ may be known, but $|\mathcal{T}\cap\mathcal{S}|$ is small; in either case simply computing a weighted average of $\mathbb{P}_{\mathcal{S}}$ and $\mathbb{P}_{\mathcal{T}}$ is not possible. Moreover, let’s also assume that the datasets consist of discrete points that are a subset of a $d=\Theta(\log n)$ dimensional boolean hypercube. The points of $\mathcal{S}$ and $\mathcal{T}$ are weighted according to probability distributions, say $\mathbb{P}_{\mathcal{S}}$ and $\mathbb{P}_{\mathcal{T}}$ respectively. Our goal is to train a neural network classifier using the labeled dataset $\mathcal{S}$ with a reweighed distribution $\mathbb{P}_{\mathcal{S}}^{\prime}$ (instead of using $\mathbb{P}_{\mathcal{S}}$), so that for training using the basic Stochastic Gradient Descent (SGD) (Robbins and Monro, 1951) with mean squared error loss, the limiting distribution [^6] In the limit as training time goes to $\infty$ and step size goes to 00. of neural network weight parameters is closer (in $\ell_{1}$ distance) to the one that would be obtained had we trained using $\mathcal{T}$ and $\mathbb{P}_{\mathcal{T}}$.
Furthermore, in general, one may not want to reweigh $\mathbb{P}_{\mathcal{S}}$ so that the resulting trained neural net behaves as if trained on $\mathbb{P}_{\mathcal{T}}$, but only tilt $\mathbb{P}_{\mathcal{S}}$ towards $\mathbb{P}_{\mathcal{T}}$ to achieve part of that effect. For example, reweigh $\mathbb{P}_{\mathcal{S}}$ to a distribution $\mathbb{P}_{\mathcal{S}}^{\prime}$ so that the post-training weight parameters of the neural network have the limiting distribution that would have resulted from the training set weights set to $(1-\alpha)\mathbb{P}_{\mathcal{S}}+\alpha\mathbb{P}_{\mathcal{T}}$. This is the case with our drug discovery example.
The boolean hypercube assumption is critically used in the proof of Theorem 3.4 (in Lemma B.11). But, it is also worth noting here that restricting the state space to a boolean hypercube is not as limiting as it may seem. First, it is natural for some applications (like drug discovery where molecules are represented as binary fingerprints). Second, the main constraint in the boolean hypercube assumption is that the underlying metric becomes $\ell_{1}$ (as binary strings can encode most reasonable inputs after discretization). However, it is well known (starting from Bourgain (1985)) that one can embed arbitrary metrics into $\ell_{1}$ with low distortion. Thus, even if the domain is not the hypercube, one can preprocess and embed it into the hypercube (with a small loss of approximation factor) and this would be better than the baseline, i.e., not reweighting. Even without any preprocessing, if there is intuitive reason to believe that the distortion will be low (the input metric is already close to $\ell_{1}$), then it is likely worth reweighting using the greedy algorithm, than not reweighting.
3.1 Drug Discovery example
@@
To illustrate with an example, for our drug discovery application in Section 4: the labeled training set $\mathcal{S}$ consists of a subset of small molecules that are binders and non-binders for the protein MNK2, and the set $\mathcal{T}$ consists of a subset of small molecules [^7] As an aside, each small molecule is usually mapped to a 2K character long binary string (fingerprint) of features. Thus, in this context, one may think of the underlying space of small molecules as a subset of the boolean hypercube in dimension 2K. labeled non-binders (non-hits) for the protein MNK1. Here the labels of the molecules in $\mathcal{T}$ are known but not necessarily on the same molecules as $\mathcal{S}$ (since the data is collected at different times with different assays). The corresponding weight distributions $\mathbb{P}_{\mathcal{S}}$ and $\mathbb{P}_{\mathcal{T}}$ may be assumed to be uniform distributions supported on $\mathcal{S}$ and $\mathcal{T}$ respectively.
Given a new set of small molecules, one now wants to rank them so that first and foremost it is a binder for MNK2 and within that ranking we also want the non-binders for MNK1 to rank higher. Such models can allow us to make predictions on large commercially available catalogs and enrich compounds that have high likelihood to bind to MNK2 but not MNK1.
Therefore, we train a neural net using $\mathcal{S}$ to predict binders and non-binders to MNK2, but at the same time we want to take into account the binder and non-binders to MNK1. One way to accomplish this is to reweigh the labeled example points in $\mathbb{P}_{\mathcal{S}}$ using $\mathbb{P}_{\mathcal{T}}$, so that points in $\mathcal{S}$ close to those in $\mathcal{T}$ receive higher weight in the reweighed distribution, denoted $\mathbb{P}_{\mathcal{S}}^{\prime}$. Training the neural net on $\mathbb{P}_{\mathcal{S}}^{\prime}$ should then achieve our goal. However, the question is how much to change the weight of each point in $\mathcal{S}$, especially when the datasets $\mathcal{S}$ and $\mathcal{T}$ can be huge. Our algorithm below suggests one scalable approach to the problem. This formulates the drug discovery application as an instance of our problem statement. See section 4 for details.
3.2 Reweighing algorithm
The reweighing algorithm (Algorithm 1) solves the problem above of how to reweigh each point in $\mathcal{S}$ so that the trained neural net behaves as if trained on $(1-\alpha)\mathbb{P}_{\mathcal{S}}+\alpha\mathbb{P}_{\mathcal{T}}$.